# Memory-Dominated Quantum Criticality as a Universal Route to High-Temperature Superconductivity

### 09 Mar 2026 v3

@@ Abstract
$\Sigma^{R}(\omega)=\int_{0}^{\Lambda}d\lambda\frac{\rho(\lambda)}{\lambda-i\omega},$ (12)
@@
$S_{\rm eff}=\int_{\omega,\mathbf{q}}\tilde{\phi}(-\omega,-\mathbf{q})\Big[-i\omega+r+cq^{2}+\Sigma^{R}(\omega)\Big]\phi(\omega,\mathbf{q})+\cdots.$ (13)
@@
$\mathrm{Re}\,\Sigma^{R}(\omega)$ $=\int_{0}^{\Lambda}d\lambda\,\rho(\lambda)\frac{\lambda}{\lambda^{2}+\omega^{2}},$ (14)
$\mathrm{Im}\,\Sigma^{R}(\omega)$ $=\omega\int_{0}^{\Lambda}d\lambda\,\rho(\lambda)\frac{1}{\lambda^{2}+\omega^{2}}.$ (15)
@@
III.1 A. Marginal pairing from overdamped collective dynamics
Near metallic quantum critical points, electrons often couple to overdamped collective modes whose retarded propagator takes the form
$D_{R}^{-1}(\omega,\mathbf{q})=r+cq^{2}+\gamma|\omega|.$ (25)
Here $D_{R}$ denotes the effective propagator of a coarse-grained order-parameter (particle–hole) collective mode in the Hertz–Millis sense: the $|\omega|$ term is generated by the electron–hole polarization of the metallic continuum (Landau damping) [3, 4]. This object should be distinguished from the Cooper-channel polarization $\Pi(\omega)$ introduced below, which diagnoses the pairing instability itself.
The linear frequency dependence reflects Landau damping due to decay of collective excitations into the electronic continuum. After momentum integration, the pairing kernel entering the Cooper channel behaves approximately as
$D_{R}(\omega)\sim\frac{1}{|\omega|}.$ (26)
The linearized gap equation may be written schematically as
$\Delta(\omega)=\int_{T}^{\Lambda}\frac{d\omega^{\prime}}{|\omega^{\prime}|}\,g(\omega-\omega^{\prime})\,\Delta(\omega^{\prime}),$ (27)
@@
To analyze the infrared flow we eliminate fast frequency modes with $\omega^{\prime}\in[\Lambda/b,\Lambda]$, corresponding to coarse–graining of short time scales. The ladder diagram gives the leading correction
$dg\propto g^{2}\int_{\Lambda/b}^{\Lambda}\frac{d\omega^{\prime}}{\omega^{\prime}}.$ (28)
The shell integral yields
@@
Introducing the RG scale $l=\ln b$ gives
@@
The Cooper interaction therefore grows only logarithmically under infrared coarse graining. In RG language the pairing interaction is marginal. Physically, overdamped collective modes dissipate into the electronic continuum and repeatedly lose phase coherence. As a result superconducting correlations accumulate only logarithmically as the energy scale is reduced.
@@
The polarization kernel entering the Cooper channel becomes
@@
Here $\Pi(\omega)$ represents the polarization of the Cooper channel generated by the collective relaxation spectrum. It should be distinguished from the conventional bosonic polarization that appears in propagator renormalization. In the present framework $\Pi(\omega)$ directly measures the dynamical pair susceptibility produced by the slow collective modes.
@@
In the time domain this corresponds to a long–memory kernel
@@
Importantly, this singular behavior does not originate from Landau damping into a fast electronic bath. Instead it arises from the spectral accumulation of slow relaxation modes themselves.
@@
The Cooper susceptibility takes the form
@@
The effective pairing kernel therefore scales as
$g_{\mathrm{eff}}(\omega)\sim|\omega|^{-(1-\eta)}.$ (39)
Under RG rescaling $\omega\rightarrow\omega/b$ the kernel transforms as
@@
Figure 2: Infrared renormalization-group flow of the Cooper-channel interaction. Shown schematically is the flow of the dimensionless pairing coupling $g$ under infrared coarse graining. Ohmic Landau damping of collective modes produces marginal logarithmic growth, $dg/dl\sim g^{2}$, characteristic of conventional Eliashberg theory (blue). In contrast, memory–dominated dynamics generated by a flat relaxation–rate density of states yields a relevant infrared instability, $dg/dl=\alpha g$, leading to algebraic amplification of the pairing interaction (orange). The slow–mode reservoir therefore dynamically enhances intrinsic electronic pairing tendencies rather than merely renormalizing a fixed microscopic coupling.
@@
In strongly correlated materials proximate to Mott insulating phases, short–range Coulomb interactions naturally generate local singlet correlations through exchange processes. Starting from the Hubbard Hamiltonian,
@@
Within the dynamical framework developed in Sec. II, the pairing susceptibility assumes the universal tt–matrix form
$\chi_{\Delta}^{-1}(\omega)=\frac{1}{g_{\rm pair}}-\Pi(\omega),$ (48)
where $g_{\rm pair}$ characterizes the intrinsic electronic pairing scale $E_{\rm pair}$ generated by short–range correlations. Diagrammatically $\Pi(\omega)$ corresponds to the Cooper bubble, which resums particle–particle ladder diagrams.
@@
Because single–particle propagation inherits the relaxational spectral structure generated by the slow–mode reservoir, the Cooper bubble reduces to a continuum integral over relaxation rates,
@@
up to numerical factors of order unity. The finite-temperature regularization of the infrared divergence, leading to $\Pi(T)\sim 1/T$, is derived explicitly in Appendix D.
@@
In the memory–dominated regime, long–time coherence is controlled by the infrared TDOS weight $\rho_{0}$. Because the pairing amplitude itself scales as $\Delta\sim E_{\rm pair}\,\rho_{0}$, the density of phase-coherent pairs is likewise proportional to the same infrared spectral weight. Integrating out amplitude fluctuations in the presence of the slow–mode reservoir therefore yields
@@
Such long–time tails can produce 1/f1/f–type noise spectra and anomalously slow relaxation, phenomena widely reported in strange metals and cuprate superconductors. Unlike conventional overdamped criticality, where correlations decay exponentially and dynamics is effectively Markovian, the memory–dominated regime predicts non-Markovian response with persistent temporal structure. Because the same slow–mode reservoir controls both pairing enhancement and dynamical response, superconducting and normal–state anomalies become two manifestations of a single infrared spectral reorganization.
@@
Figure 4: Universality map relating the relaxation-rate spectrum to infrared collective dynamics and superconducting pairing. The TDOS exponent $\alpha$ determines the scaling of the polarization kernel $\Pi(\omega)\sim|\omega|^{\alpha-1}$. The Ohmic Hertz–Millis theory corresponds to the marginal point $\alpha=1$, while a flat TDOS ($\alpha=0$) produces memory-dominated critical dynamics with $\Pi(\omega)\sim 1/|\omega|$ and enhanced pairing.
@@
This structure is summarized schematically in Fig. 4. The figure illustrates how the infrared universality class of the collective dynamics is determined by the exponent of the relaxation-rate density of states. For a TDOS scaling $\rho(\lambda)\sim\lambda^{\alpha}$, the polarization kernel assumes the universal infrared form
$\Pi(\omega)\sim|\omega|^{\alpha-1}.$ (65)
Different dynamical regimes therefore emerge as special cases of the same spectral organization. For $\alpha>1$ the density of slow modes vanishes in the infrared, so that $\Pi(\omega)\rightarrow 0$ and the collective dynamics reduces to weakly damped propagating modes with $\chi^{-1}\simeq r+cq^{2}$. The marginal case $\alpha=1$ corresponds to Ohmic dissipation, reproducing the Hertz–Millis form $\chi^{-1}\simeq r+cq^{2}+\gamma|\omega|$. Most importantly, when the TDOS approaches a finite constant ($\alpha=0$), a dense continuum of near-marginal relaxation modes emerges. In this regime the polarization acquires the singular form $\Pi(\omega)\sim 1/|\omega|$, producing long-time memory kernels $K(t)\sim 1/t$ and the non-Markovian infrared dynamics discussed throughout this work.
The lower panel of Fig. 4 illustrates the direct consequence of this spectral organization for superconductivity. Within the Thouless criterion $\chi_{\Delta}^{-1}(0,T)=g_{\rm pair}^{-1}-\Pi(T)$, Ohmic dynamics leads to the familiar logarithmic kernel $\Pi(T)\sim\ln(\Omega/T)$ and the exponentially small BCS transition temperature. In contrast, memory-dominated dynamics with a flat TDOS generates the stronger scaling $\Pi(T)\sim C/T$, converting moderate microscopic pairing into an algebraically enhanced instability $T_{c}\sim\rho_{0}E_{\rm pair}$. The figure therefore provides a compact universality map linking the relaxation-rate spectrum, infrared collective dynamics, and the resulting pairing scale.
A direct physical consequence of this spectral reorganization is the enhancement of superconducting instabilities. Superconductivity need not rely on finely tuned bosonic mediators. Instead, intrinsic short-range electronic pairing tendencies are dynamically amplified by the slow-mode reservoir. The exponential suppression characteristic of weak-coupling BCS theory is replaced by algebraic scaling controlled by infrared spectral weight. In this sense, superconducting order corresponds to condensation within an extended slow-mode manifold rather than binding via a sharp bosonic excitation.
@@
The relaxation-rate organization emphasized here is not merely a phenomenological assumption. Both Liouvillian evolution in open quantum systems and generalized memory-kernel dynamics naturally decompose into continua of decay modes characterized by complex relaxation eigenvalues. Whenever long-time correlations are present, Laplace-domain representations generically produce continuous relaxation spectra. A dense TDOS therefore reflects a universal mathematical property of dynamical systems with extended temporal correlations rather than a model-specific constructione, is derived in Appendix F.
@@
In contrast, the relaxation-spectrum formulation adopted here emphasizes the infrared organization of time scales themselves. When the density of relaxation rates develops a finite limit $\rho(\lambda\to 0)=\rho_{0}$, the system possesses a continuum of slow modes extending to arbitrarily long time scales. The resulting dynamics naturally exhibits scale-free temporal behavior, including long-time memory kernels $K(t)\sim 1/t$ and logarithmic infrared response functions. Scale-free temporal dynamics therefore arises directly from the structure of the relaxation spectrum rather than secondarily from spatial criticality. This perspective is closely related to the scale-free lifetime statistics observed in self-organized critical systems and suggests that infrared relaxation spectra provide a natural language for describing dynamical critical phenomena.
Memory–dominated dynamics as a guiding principle for room-temperature superconductivity. Most contemporary efforts toward room-temperature superconductivity focus on strengthening microscopic pairing interactions through pressure, lattice engineering, or exotic chemical composition. Within the present framework, however, enhancing the bare pairing scale alone is insufficient. The decisive control parameter for high transition temperatures is the infrared dynamical organization of collective modes encoded in the relaxation-rate density of states.
@@
The present theory therefore suggests a complementary design principle for superconductivity beyond material-specific glue engineering: stabilization of dynamical phases that support extensive slow-mode continua [27]. Control of electronic frustration, dimensionality, feedback mechanisms, or proximity to dynamical criticality may be as crucial as increasing microscopic attraction itself. In this perspective high-temperature superconductivity corresponds to a dynamical phase characterized by maximal infrared memory rather than simply by extreme coupling strength.
@@
In this perspective, high–temperature superconductivity appears as a generic consequence of infrared spectral organization in strongly correlated matter. Systematic exploration of TDOS-controlled dynamics may provide a pathway toward stabilizing and enhancing collective quantum states in future materials and devices.

### 16 Mar 2026 v4

@@ Abstract
$\Sigma_{R}(\omega)=\int_{0}^{\Lambda}d\lambda\frac{\rho(\lambda)}{\lambda-i\omega},$ (12)
@@
$S_{\rm eff}=\int_{\omega,\mathbf{q}}\tilde{\phi}(-\omega,-\mathbf{q})\Big[-i\omega+r+cq^{2}+\Sigma_{R}(\omega)\Big]\phi(\omega,\mathbf{q})+\cdots.$ (13)
@@
$\mathrm{Re}\,\Sigma_{R}(\omega)$ $=\int_{0}^{\Lambda}d\lambda\,\rho(\lambda)\frac{\lambda}{\lambda^{2}+\omega^{2}},$ (14)
$\mathrm{Im}\,\Sigma_{R}(\omega)$ $=\omega\int_{0}^{\Lambda}d\lambda\,\rho(\lambda)\frac{1}{\lambda^{2}+\omega^{2}}.$ (15)
@@
III.1 A. Marginal pairing in overdamped bosonic-mediator theories
As a reference point, it is useful to recall the structure of pairing in conventional theories based on overdamped collective mediators, such as spin-fluctuation approaches near metallic quantum criticality [25, 26]. In such treatments one first introduces an effective bosonic propagator for a particle–hole collective mode,
$D_{R}^{-1}(\omega,\mathbf{q})=r+cq^{2}+\gamma|\omega|,$ (25)
where the $|\omega|$ term arises from Landau damping by the electronic continuum. This object describes the retarded susceptibility of an overdamped order-parameter fluctuation and should not be identified with the Cooper-channel polarization kernel itself.
In bosonic-mediator theories [27], the overdamped collective mode is then treated as an effective pairing mediator. After momentum integration, the resulting interaction kernel entering the Cooper channel is schematically of the form
$V(\omega)\sim\frac{1}{|\omega|},$ (26)
up to model-dependent form factors and ultraviolet regularization. If one further assumes that pairing can still be analyzed within a quasiparticle-based ladder or gap-equation framework, the linearized pairing equation takes the schematic form
$\Delta(\omega)=\int_{T}^{\Lambda}\frac{d\omega^{\prime}}{|\omega^{\prime}|}\,V(\omega-\omega^{\prime})\,\Delta(\omega^{\prime}),$ (27)
@@
The corresponding infrared flow can be understood by integrating out fast frequency modes in the shell $\omega^{\prime}\in[\Lambda/b,\Lambda]$. The leading ladder correction is then
$dg\propto g^{2}\int_{\Lambda/b}^{\Lambda}\frac{d\omega^{\prime}}{\omega^{\prime}},$ (28)
which yields
@@
Introducing the RG scale $l=\ln b$, one obtains the standard marginal flow
@@
Thus, within overdamped bosonic-mediator theories, pairing grows only logarithmically in the infrared [28, 29]. The physical reason is that the mediator itself is continuously damped by decay into the electronic continuum, so coherence is not retained across an extended hierarchy of time scales. The resulting enhancement is therefore marginal rather than algebraic.
This benchmark should be distinguished sharply from the framework developed in the present work. Here we do not assume a well-defined quasiparticle description nor an effective pairing glue represented by a single bosonic propagator. Instead, after integrating out microscopic electronic degrees of freedom, the infrared Cooper sector is organized by the collective relaxation spectrum itself. The relevant object is not the retarded propagator of an overdamped order-parameter mode, but the polarization kernel generated by a continuum of relaxational pair modes, as derived below.
@@
While the quantity above governs the retarded response of the collective sector, the object relevant for superconducting pairing is the Cooper-channel fluctuation kernel. After integrating out microscopic electronic degrees of freedom, pairing correlations are encoded in the two-point correlator of the collective pair field
$\langle\Delta(\omega)\Delta^{*}(\omega)\rangle.$ (32)
To understand its structure, consider a single relaxational pair mode with decay rate $\lambda$. At long times its dynamics is well approximated by a linear stochastic equation
$\dot{\Delta}_{\lambda}=-\lambda\,\Delta_{\lambda}+\xi_{\lambda},$ (33)
where $\xi_{\lambda}$ denotes Gaussian noise generated by microscopic degrees of freedom. The stationary two-point correlation function of such a mode is
$\langle\Delta_{\lambda}(t)\Delta_{\lambda}^{*}(0)\rangle\propto e^{-\lambda|t|}.$ (34)
Fourier transforming this correlation function gives the fluctuation spectrum of a relaxational mode
$S_{\lambda}(\omega)\propto\frac{1}{\lambda^{2}+\omega^{2}}.$ (35)
The Cooper-channel polarization kernel therefore arises by summing the contributions of all relaxational pair modes weighted by the TDOS,
@@
Equation (36) represents the dynamical polarization of the Cooper channel generated by the collective relaxation spectrum. Importantly, this object differs from the reactive part of the retarded susceptibility discussed above. Rather than describing the propagation of a single bosonic mediator, it measures the fluctuation spectrum of the entire slow-mode reservoir.
@@
In the time domain this corresponds to a long-memory kernel
@@
Importantly, this singular behavior does not originate from Landau damping into a fast electronic bath. Instead it arises from the spectral accumulation of slow relaxation modes themselves. A complete derivation of Eq. (36) from the stochastic dynamics of relaxational pair modes is presented in Appendix D.
@@
In conventional superconductivity theories the transition temperature is often obtained from a microscopic gap equation derived from fermionic quasiparticle propagators. Such formulations rely on the existence of reasonably well-defined single-particle excitations, so that the Cooper kernel can be written in terms of products of fermionic Green functions.
In strongly correlated systems, however, the quasiparticle picture may cease to be valid. The single-particle spectral function can become broad and incoherent, and the low-energy dynamics may instead be governed by a continuum of collective relaxation processes. Within the present framework the infrared dynamics is therefore organized in terms of the relaxation-rate spectrum $\rho(\lambda)$ rather than quasiparticle propagation.
Because the microscopic quasiparticle propagator is not introduced explicitly, the superconducting transition is identified through the instability of the Cooper-channel susceptibility. The corresponding Thouless criterion reads
$\chi_{\Delta}^{-1}(0,T_{c})=0,\quad\text{or equivalently}\quad 1-g_{\rm pair}\Pi(T_{c})=0.$ (40)
Here $g_{\rm pair}$ represents the intrinsic electronic interaction that seeds Cooper correlations, while the relaxation spectrum controls their dynamical amplification through the infrared structure of $\Pi(\omega)$.
The Cooper susceptibility therefore takes the form
@@
The Cooper-channel kernel therefore acquires the infrared scaling
$\Pi(\omega)\sim|\omega|^{\eta-1}.$ (44)
Under RG rescaling $\omega\rightarrow\omega/b$ the effective pairing interaction transforms as
@@
Figure 2: Infrared renormalization-group flow of the Cooper-channel interaction. Shown schematically is the flow of the dimensionless pairing coupling $g$ under infrared coarse graining. Ohmic Landau damping of collective modes produces marginal logarithmic growth, $dg/dl\sim g^{2}$, characteristic of conventional Eliashberg theory (blue). In contrast, memory–dominated dynamics generated by a flat relaxation–rate density of states yields a relevant infrared instability, $dg/dl=(1-\eta)g$, leading to algebraic amplification of the pairing interaction (orange). The slow–mode reservoir therefore dynamically enhances intrinsic electronic pairing tendencies rather than merely renormalizing a fixed microscopic coupling.
@@
In strongly correlated materials proximate to Mott insulating phases, short–range Coulomb interactions naturally generate local singlet correlations through exchange processes [1, 30, 31]. Starting from the Hubbard Hamiltonian,
@@
Within the dynamical framework developed in Sec. III, the pairing susceptibility assumes the universal tt–matrix form in Eq. (41), where $g_{\rm pair}$ characterizes the intrinsic electronic pairing scale $E_{\rm pair}$ generated by short–range correlations. Diagrammatically this corresponds to the Cooper-channel polarization obtained after integrating out microscopic electronic degrees of freedom.
@@
Because the collective pair fluctuations inherit the relaxational spectral structure generated by the slow-mode reservoir, the Cooper-channel polarization kernel reduces to a continuum integral over relaxation rates,
@@
At finite temperature the imaginary-time direction becomes compact, $\tau\in[0,\beta]$ with $\beta=1/T$. As a consequence the system cannot probe arbitrarily long time scales. In the imaginary-time formalism the dynamical fields are expanded in bosonic Matsubara frequencies
$\Omega_{n}=2\pi nT.$ (56)
The smallest nonzero dynamical frequency therefore sets a natural infrared scale of order
$|\omega|\sim T.$ (57)
Substituting this scale into the infrared expression (55) gives
$\Pi(T)\sim\frac{\rho_{0}}{T}.$ (58)
Thus the infrared divergence present at zero temperature is universally regulated by the thermal scale.
@@
up to numerical factors of order unity.
@@
In the memory–dominated regime, long–time coherence is controlled by the infrared TDOS weight $\rho_{0}$. Because the pairing amplitude itself scales as $\Delta\sim E_{\rm pair}\,\rho_{0}$, the density of phase-coherent pairs is likewise proportional to the same infrared spectral weight. Integrating out amplitude fluctuations in the presence of the slow-mode reservoir generically yields
@@
Such long-time tails naturally produce 1/f-type noise spectra and anomalously slow relaxation, phenomena widely reported in strange metals and cuprate superconductors. Unlike conventional overdamped criticality, where correlations decay exponentially and dynamics is effectively Markovian, the memory–dominated regime predicts non-Markovian response with persistent temporal structure. Because the same slow–mode reservoir controls both pairing enhancement and dynamical response, superconducting and normal–state anomalies become two manifestations of a single infrared spectral reorganization.
@@
Figure 4: Dynamical classification determined by the TDOS exponent $\alpha$. The scaling $\rho(\lambda)\sim\lambda^{\alpha}$ implies $\Pi(\omega)\sim|\omega|^{\alpha-1}$ and separates the system into memory–dominated ($\alpha<1$), marginal Ohmic ($\alpha=1$), and analytic ($\alpha>1$) regimes, with corresponding pairing consequences. Here $\omega$ denotes Matsubara frequency.
@@
This structure is summarized schematically in Fig. 4. The figure illustrates how the infrared universality class of the collective dynamics is determined by the exponent of the relaxation-rate density of states. For a TDOS scaling $\rho(\lambda)\sim\lambda^{\alpha}$, the polarization kernel assumes the universal infrared form $\Pi(\omega)\sim|\omega|^{\alpha-1}$. Different dynamical regimes therefore emerge as special cases of the same spectral organization. For $\alpha>1$ the density of slow modes vanishes in the infrared, so that the polarization kernel remains finite and the collective dynamics approaches the weakly damped limit, with the infrared response dominated by the nondissipative part $\chi^{-1}\simeq r+cq^{2}$. The marginal case $\alpha=1$ corresponds to Ohmic dissipation, reproducing the Hertz–Millis form $\chi^{-1}\simeq r+cq^{2}+\gamma|\omega|$. Most importantly, when the TDOS approaches a finite constant ($\alpha=0$), a dense continuum of near-marginal relaxation modes emerges. In this regime the polarization acquires the singular form $\Pi(\omega)\sim 1/|\omega|$, producing long-time memory kernels $K(t)\sim 1/t$ and the non-Markovian infrared dynamics discussed throughout this work.
Interestingly, the infrared response generated by a flat TDOS is structurally identical to anomalous dynamical susceptibilities that have appeared in several apparently unrelated contexts [5, 32]. In particular, the resulting spectrum corresponds in frequency space to a logarithmic real part and a nearly constant absorptive component of the susceptibility. Similar infrared structures were introduced phenomenologically in the marginal Fermi liquid theory of strange metals and also arise in solvable strongly correlated spin models such as the Sachdev–Ye spin liquid. In the present framework, however, this behavior follows directly from the accumulation of slow relaxation modes and therefore emerges naturally as a consequence of a finite infrared relaxation-rate density of states.
The lower panel of Fig. 4 illustrates the direct consequence of this spectral organization for superconductivity. Within the Thouless criterion $\chi_{\Delta}^{-1}(0,T)=g_{\rm pair}^{-1}-\Pi(T)$, Ohmic dynamics leads to the familiar logarithmic Cooper kernel $\Pi(T)\sim\ln(\Omega/T)$ and hence to the exponentially small transition scale characteristic of conventional marginal pairing theories. In contrast, memory-dominated dynamics with a flat TDOS generates the stronger scaling $\Pi(T)\sim C/T$, converting moderate microscopic pairing into an algebraically enhanced instability $T_{c}\sim\rho_{0}E_{\rm pair}$. The figure therefore provides a compact universality map linking the relaxation-rate spectrum, infrared collective dynamics, and the resulting pairing scale.
A direct physical consequence of this spectral reorganization is the enhancement of superconducting instabilities. Superconductivity need not rely on finely tuned bosonic mediators. Instead, intrinsic short-range electronic pairing tendencies are dynamically amplified by the slow-mode reservoir. The exponential suppression characteristic of weak-coupling BCS theory is replaced by algebraic scaling controlled by infrared spectral weight. In this sense, superconducting order is governed by condensation in a Cooper sector embedded within an extended slow-mode manifold rather than by binding through a sharp bosonic excitation.
@@
The relaxation-rate organization emphasized here is not merely a phenomenological assumption. Both Liouvillian evolution in open quantum systems and generalized memory-kernel dynamics naturally decompose into continua of decay modes characterized by complex relaxation eigenvalues. Whenever long-time correlations are present, Laplace-domain representations generically produce continuous relaxation spectra. A dense TDOS therefore reflects a generic mathematical feature of dynamical systems with extended temporal correlations rather than a model-specific construction. Related derivations are presented in Appendix F.
@@
In contrast, the relaxation-spectrum formulation adopted here emphasizes the infrared organization of time scales themselves. When the density of relaxation rates develops a finite limit $\rho(\lambda\to 0)=\rho_{0}$, the system possesses a continuum of slow modes extending to arbitrarily long time scales. The resulting dynamics naturally exhibits scale-free temporal behavior, including long-time memory kernels $K(t)\sim 1/t$ and nonanalytic infrared response functions, with logarithmic or stronger singular structures depending on the observable sector. Scale-free temporal dynamics therefore arises directly from the structure of the relaxation spectrum rather than secondarily from spatial criticality. This perspective is closely related to the scale-free lifetime statistics observed in self-organized critical systems and suggests that infrared relaxation spectra provide a natural language for describing dynamical critical phenomena [33, 34].
Memory–dominated dynamics as a guiding principle for room-temperature superconductivity. Most contemporary efforts toward room-temperature superconductivity focus on strengthening microscopic pairing interactions through pressure, lattice engineering, or exotic chemical composition. Within the present framework, however, enhancing the bare pairing scale alone is insufficient. A decisive control parameter for high transition temperatures may be the infrared dynamical organization of collective modes encoded in the relaxation-rate density of states.
@@
The present theory therefore suggests a complementary design principle for superconductivity beyond material-specific glue engineering: stabilization of dynamical phases that support extensive slow-mode continua [35]. Control of electronic frustration, dimensionality, feedback mechanisms, or proximity to dynamical criticality may be as crucial as increasing microscopic attraction itself. In this perspective high-temperature superconductivity is most naturally favored in regimes characterized by maximal infrared memory rather than simply by extreme coupling strength.
@@
In this perspective, high–temperature superconductivity can be viewed as a generic consequence of favorable infrared spectral organization in strongly correlated matter. Systematic exploration of TDOS-controlled dynamics may provide a pathway toward stabilizing and enhancing collective quantum states in future materials and devices.